# Visual Attention Analysis in Online Learning

### v2

@@ Abstract
The second one, M2LADS, facilitates the integration and visualization of multimodal data captured during learning sessions within a MOOC, presented through Web-based Dashboards.
@@
With the main aim to complement and provide analysis of some data managed by M2LADS, we have developed a tool designed to visualize eye movement data collected during a learning session (LS) in online learning. The name of this tool is VAAD, which stands for Visual Attention Analysis Dashboard.
In this article we present the tool VAAD. The structure of this article is as follows: In the following section we present related works and the motivation for the approach proposed. Then, in Section III we provide a description of the context and dataset. In Section IV, a detailed explanation of the proposed approach is presented. Finally, in Section V the article concludes with conclusions and future work.
@@ III Context and Dataset
In order to test the approach proposed, we have conducted a study in the School of Engineering at our institution, where 120 learners from the school were monitored in our laboratory while they attended and interacted with a MOOC subunit during a 30-minute LS. The chosen MOOC was titled ”Introduction to Development of Web Application” (WebApp), which is available on the edX MOOC platform111https://www.edx.org/ and offered by our university.
@@ IV Approach proposed
The VAAD tool is composed of three different modules: a processing data module, a visualization module, and a prediction module. As seen in Fig. 1, the processing module manages the data from the eye-tracker and M2LADS, and is interconnected with both the visualization and prediction modules, which use the processed data for their respective tasks.
@@ IV-A Processing Data Module
Before the processing stage, the eye-tracking data of learners were thoroughly inspected to identify errors or anomalies. Learners displaying a notable frequency of ”Eyes Not Found” values for the Eye movement type variable were excluded to ensure the integrity of the final data utilized by VAAD.
@@ IV-B Visualization Module
Another interactive chart presents individual learners’ data through a heat map, providing a visual representation of their on-screen attention. As depicted in the example shown in Fig. 3, it is possible to identify precisely where the learner was looking during the video task, with a higher concentration of fixations observed within the region of the screen where the video was displayed. Analysts can select a specific learner and customize their visual screen heat map by choosing from various options, such as different activities within the MOOC session. This approach offers valuable insights into each learner’s focus during different activities, assisting in the identification of engaging materials and assessing the impact of diverse learning sources on learners’ performance.
@@ V Conclusions and Future work
In this paper, we introduce VAAD (Visual Attention Analysis Dashboard), an innovative tool designed to visualize biometric data related to visual attention gathered from an online learning session. This tool enables analysts to gain a deeper understanding of learner behavior by filtering and visualizing different components of the session.
@@
Finally, future work will further explore current and emerging predictive methods to ascertain the most suitable model for task prediction and investigate other indicators that have proven to be useful in predicting activities such as eyeblink [29, 30], eye pupil size [31], keystroking [32], among others.
@@ Acknowledgment
Support by projects: HumanCAIC (TED2021-131787B-I00 MICINN), SNOLA (RED2022-134284-T), e-Madrid-CM (S2018/TCS-4307, a project which is co-funded by the European Structural Funds, FSE and FEDER), IndiGo! (PID2019-105951RB-I00), TEA360 (PID2023-150488OB-I00, SPID202300X150488IV0) and BIO-PROCTORING (GNOSS Program, Agreement Ministerio de Defensa-UAM-FUAM dated 29-03-2022).Roberto Daza is supported by a FPI fellowship from MINECO/FEDER.

### v4

# VAAD: Visual Attention Analysis Dashboard applied to e-Learning

@@ Abstract
The second one, M2LADS, facilitates the integration and visualization of multimodal data captured during monitoring using biometric sensors, presented through Web-based Dashboards.
With the main aim to complement and provide analysis of data managed by e-learning platforms, we have developed a tool designed to visualize eye movement data collected. The name of this tool is VAAD, which stands for Visual Attention Analysis Dashboard. This tool presents visualizations through different graphs such as boxplots and heatmaps to help understand the monitored individuals’ behavior. In this article, we use data gathered during a learning session (LS) in online learning to better understand learners’ behavior. However, VAAD could also be applied in other contexts involving web platforms and services, such as marketing studies or system analysis [11], to gain insights into users’ preferences, or in gaming to improve players’ performance.
@@
In this article we present the tool VAAD applied to e-learning. The structure of this article is as follows: In the following section we present related works and the motivation for the approach proposed. Then, in Section III we provide a description of the context and dataset. In Section IV, a detailed explanation of the proposed approach is presented. Finally, in Section V the article concludes with conclusions and future work.
@@ III Context and Dataset
In order to test the approach proposed, we have conducted a study in the School of Engineering at our institution, where 120 learners from the school were monitored in our laboratory while they attended and interacted with a MOOC subunit during a 30-minute LS. The chosen MOOC was titled "Introduction to Development of Web Application" (WebApp), which is available on the edX MOOC platform111https://www.edx.org/ and offered by our university.
@@ IV Approach proposed
The VAAD tool is composed of three different modules: a processing data module, a visualization module, and a prediction module. As seen in Fig. 1, the processing module manages the data from the eye-tracker and M2LADS, and is interconnected with both the visualization and prediction modules, which use the processed data for their respective tasks. Finally, all preprocessed data and results are stored in M2LADS database.
@@ IV-A Processing Data Module
Before the processing stage, the eye-tracking data of learners were thoroughly inspected to identify errors or anomalies. Learners displaying a notable frequency of "Eyes Not Found" values for the Eye movement type variable were excluded to ensure the integrity of the final data utilized by VAAD.
@@ IV-B Visualization Module
Another interactive chart presents individual learners’ data through a heat map, providing a visual representation of their on-screen attention. As depicted in the example shown in Fig. 3 (a), it is possible to identify precisely where the learner was looking during the video task, with a higher concentration of fixations observed within the region of the screen where the video was displayed (Fig. 3 (b)). Analysts can select a specific learner and customize their visual screen heat map by choosing from various options, such as different activities within the MOOC session. This approach offers valuable insights into each learner’s focus during different activities, assisting in the identification of engaging materials and assessing the impact of diverse learning sources on learners’ performance.
@@ V Conclusions and Future work
In this paper, we introduce VAAD (Visual Attention Analysis Dashboard), an innovative tool designed to visualize biometric data related to visual attention gathered from monitored sessions. In this article, we use eye movement data from a learning session and this tool enables analysts to gain a deeper understanding of learner behavior by filtering and visualizing different components of the session.
@@
Finally, future work will further explore current and emerging predictive methods to ascertain the most suitable model for task prediction and investigate other indicators that have proven to be useful in predicting activities such as eyeblink [30, 31], eye pupil size [32], keystroking [33], among others. Additionally, we will tested the use of this tool in others contexts where the behavior of users looking at the screen is to be analyzed.
@@ Acknowledgment
Support by projects: HumanCAIC (TED2021-131787B-I00 MICINN), SNOLA (RED2022-134284-T), e-Madrid-CM (S2018/TCS-4307, a project which is co-funded by the European Structural Funds, FSE and FEDER), IndiGo! (PID2019-105951RB-I00), TEA360 (PID2023-150488OB-I00, SPID202300X150488IV0), BIO-PROCTORING (GNOSS Program, Agreement Ministerio de Defensa-UAM-FUAM dated 29-03-2022), and Cátedra ENIA UAM-VERIDAS en IA Responsable (NextGenerationEU PRTR TSI-100927-2023-2). Roberto Daza is supported by a FPI fellowship from MINECO/FEDER.